Farmer Jessica's Produce Stand sells 21 kg bags of mixed nuts that contain 38% peanuts. To make her product she combines Brand A mixed nuts which contain 40% peanuts and Brand B mixed nuts which contain 35% peanuts. How much of each does she need to use?
Answer by checkley77(12844)

.40X+.35(21-X)=.38*21
.40X+7.35-.35X=7.98
.05X=7.98-7.35
.05X=.63
X=63/.05
X=12.6 KG OF THE 40% MIX IS USED.
21-12.6=8.4 KG OR THE 35% MIX IS USED.
PROOF:
.40*12.6+.35*8.4=7.98
5.04+2.94=7.98
7.98=7.98